OpenStudy (anonymous):

Hey Can you help with some math? ? @loser66

OpenStudy (anonymous):

i can help if you want

OpenStudy (anonymous):

Yes thank you

OpenStudy (anonymous):

What is your question?

OpenStudy (anonymous):

Theres lots of parts, so ill post them one part at a time

OpenStudy (anonymous):

Part A The first part of Ray and Kelsey's roller coaster is a curved pattern that can be represented by a polynomial function. Ray and Kelsey are working to graph a third-degree polynomial function that represents the first pattern in the coaster plan. Ray says the third-degree polynomial has 4 intercepts. Kelsey argues the function can have as many as 3 zeros only. Is there a way for the both of them to be correct? Explain your answer. Kelsey has a list of possible functions. Pick one of the g(x) functions below and then describe to Kelsey the key features of g(x), including the end behavior, y-intercept, and zeros. g(x) = x3 − x2 − 4x + 4 g(x) = x3 + 2x2 − 9x − 18 g(x) = x3 − 3x2 − 4x + 12 g(x) = x3 + 2x2 − 25x − 50 g(x) = 2x3 + 14x2 − 2x − 14 Create a graph of the polynomial function you selected from Question 2.

OpenStudy (anonymous):

hold up im doin it

OpenStudy (anonymous):

Ookay

OpenStudy (anonymous):

I apologize I cannot help you with this I was never good at polynomial. But @jcoury will probably help you.

OpenStudy (anonymous):

Its okay, thankyou

OpenStudy (anonymous):

a third degree polynomial can have at most three x intercepts and always has one y intercept. So it can have 4 intercepts. And three zeroes . they are both correct

OpenStudy (anonymous):

So thats what i put for he first part of Ray and Kelsey's roller coaster is a curved pattern that can be represented by a polynomial function. Ray and Kelsey are working to graph a third-degree polynomial function that represents the first pattern in the coaster plan. Ray says the third-degree polynomial has 4 intercepts. Kelsey argues the function can have as many as 3 zeros only. Is there a way for the both of them to be correct? Explain your answer. ?

OpenStudy (anonymous):

able to cross the x three time and the y once or depending where the exponents are either on x or y it can cross the y three times and the x once if you are looking for it crossing any and all axis (x and y) then they cross 4 times if a specific axis then it depends on the axis in shorter words the girl is only looking at one axis and the boy all

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

they are both right but im trying to find out what graph

OpenStudy (anonymous):

They didnt give us any graphs

OpenStudy (anonymous):

thats not what i mean basically pick an equation and create a corresponding graph

OpenStudy (anonymous):

ooh idk :\

OpenStudy (anonymous):

okay pick an equation. When you picked one tell me.

OpenStudy (anonymous):

g(x) = x3 − x2 − 4x + 4 the first one?

OpenStudy (anonymous):

Btw could you fan me?

OpenStudy (anonymous):

you may want to pick the second graph though

OpenStudy (anonymous):

Why?

OpenStudy (anonymous):

i dont get any of this lol